# Robustness of neutron star merger simulations to changes in neutrino transport and neutrino-matter interactions

### v1

@@ III.3 Neutrinos in the remnant
Finally, on Fig. 11 we report the luminosity of neutrinos in the $(1-3)\,{\rm ms}$ post-merger for which data is available for all simulations. The most noticeable feature here is obviously the large burst of $\bar{\nu}_{e}$ emission from simulation M1-SpEC over a $0.5\,{\rm ms}$ period following the merger. We already noted that the high-density regions of the remnant erroneously neutronize in this simulation; this is another indication that the early post-merger composition evolution of the high-density regions is not well captured in M1-SpEC. A more detailed analysis of the neutrino distribution in post-processing shows that during a $\sim 0.5\,{\rm ms}$ period around merger, the M1-SpEC simulation created large numbers of low-energy $\bar{\nu}_{e}$ in the dense, cold regions of the merging neutron star cores, and that lepton number conservation was not satisfied during that time period (as opposed to what is observed in other simulations using the various M1 schemes). These low-energy $\bar{\nu}_{e}$ are the source of the extra emission of $\bar{\nu}_{e}$ at early times, but also of the decrease in the electron fraction observed in dense regions, as the conservation of lepton number was satisfied during their partial reabsorption into the remnant. The exact source of this violation of lepton number conservation is unclear, but likely due to a combination of the inaccuracies of the M1-SpEC scheme in dense cold regions and of corrections applied to the evolution of the number density to guarantee that the average energy of neutrinos remain between $0.01\,{\rm MeV}$ and $1000\,{\rm MeV}$. This is of course a clear indication that the results of the other simulations should be trusted over the M1-SpEC simulation in any region impacted by this problem.
@@ IV Discussion
Our simulations also allow us to assess more directly the impact of improvements to the M1 scheme recently proposed by Radice et al [39]. We find indications that these improvements do indeed modify the behavior of neutrinos in regions with high scattering optical depths, as predicted in [39], and generally lead to differences with our preexisting M1 code that are often comparable to or larger than other sources of uncertainties in the neutrino transport algorithm. We note in addition that the M1-SpEC simulation has a peculiar behavior leading to the emission of low-energy $\bar{\nu}_{e}$ during merger, violating conservation of electron lepton number. This effect has not been seen in other M1 simulations, yet considering that it occurs in the same high density, cold regions where the M1-SpEC is expected to be least accurate, it is likely a result of the inaccuracies of that M1 scheme.
The M1-Radice and MC simulations show $\sim 20\%$ level agreements in early-time neutrino luminosities, while more significant differences are observed between those three simulations and those using the approximate implicit solve previously implemented in SpEC. Interestingly, this difference between M1-Radice and M1-SpEC also appears sufficient to explain the largest difference between M1 and MC results observed in our previous comparison of low-mass binary neutron star systems, i.e. the fact that the M1 scheme was underproducing $\nu_{x}$ neutrinos by up to $\sim 50\%$.

### v2

@@ III.3 Neutrinos in the remnant
On Fig. 11 we report the luminosity of neutrinos in the $(1-3)\,{\rm ms}$ post-merger for which data is available for all simulations. The most noticeable feature here is obviously the large burst of $\bar{\nu}_{e}$ emission from simulation M1-SpEC over a $0.5\,{\rm ms}$ period following the merger. We already noted that the high-density regions of the remnant erroneously neutronize in this simulation; this is another indication that the early post-merger composition evolution of the high-density regions is not well captured in M1-SpEC. A more detailed analysis of the neutrino distribution in post-processing shows that during a $\sim 0.5\,{\rm ms}$ period around merger, the M1-SpEC simulation created large numbers of low-energy $\bar{\nu}_{e}$ in the dense, cold regions of the merging neutron star cores, and that lepton number conservation was not satisfied during that time period. These low-energy $\bar{\nu}_{e}$ are the source of the extra emission of $\bar{\nu}_{e}$ at early times, but also of the decrease in the electron fraction observed in dense regions, as the conservation of lepton number was satisfied during their partial reabsorption into the remnant. The exact source of this violation of lepton number conservation is unclear, but likely due to a combination of the inaccuracies of the M1-SpEC scheme in dense cold regions and of corrections applied to the evolution of the number density to guarantee that the average energy of neutrinos remain between $0.01\,{\rm MeV}$ and $1000\,{\rm MeV}$. This is of course a clear indication that the results of the other simulations should be trusted over the M1-SpEC simulation in any region impacted by this problem.
@@
This agreement between M1 and MC results on the total luminosity does not imply that the distributions of neutrinos in low-density regions are similar. In Fig. 12, we show the distribution of neutrinos with respect to the angle between the neutrino momentum (or, for M1, flux) and the polar axis. We see that the M1 simulation has significantly more neutrinos moving in the polar direction, and significantly less in the equatorial direction, especially for $\nu_{x}$. This is a known consequence of the use of the approximate Minerbo closure, which results in neutrinos coming from different regions of the remnant colliding and forming a beam of neutrinos in the polar direction. Such an effect has also been seen for different M1 implementations in e.g. [27, 73].
Figure 12: Distribution of neutrinos with respect to polar angle of propagation $\theta$, for each species of neutrinos. We show results for the M1-Radice (thick lines) and MC-HR (thin lines) simulations, for neutrinos leaving the computational domain during the $\sim 3\,{\rm ms}$ preceding the collapse of the neutron star in MC-HR. We measure the angle between the neutrino momentum and the z-axis for MC, and between the flux vector and the z-axis for M1. The M1 simulation has significantly larger neutrino density in the polar direction, and significantly smaller neutrino density in the equatorial direction.
Sim $N_{l}/N_{l}^{0}$ $N_{l,fluid}/N_{l}^{0}$ $N_{l,\nu}/N_{l}^{0}$ $N_{l,outflow}/N_{l}^{0}$
MC-HR 1.003 1.149 -0.146 -0.009
M1-Radice 0.985 1.089 -0.104 -0.015
M1-NuLib 1.010 1.120 -0.110 -0.015
M1-SpEC 0.677 0.932 -0.255 -0.289
Table 4: Net lepton number on the grid $N_{l}=N_{e^{-}}-N_{e^{+}}+N_{\nu_{e}}-N_{\bar{\nu}_{e}}$ $2.5\,{\rm ms}$ post-merger, normalized by its initial value $N_{l}^{0}$. We also show the net lepton number of the fluid (electrons and positrons) and neutrinos ($\nu_{e}$ and $\bar{\nu}_{e}$) on the grid, and of the neutrinos that left the grid. For exact conservation of lepton number, the second column plus the last column should be exactly 1.
Finally, we can look more carefully at lepton number evolution and conservation in the simulations, using information from the neutrino sector. In Table 4, we show the total lepton number $N_{l}=N_{e^{-}}-N_{e^{+}}+N_{\nu_{e}}-N_{\bar{\nu}_{e}}$ in all simulations $2.5\,{\rm ms}$ post-merger, normalized by its initial value. We also show how that lepton number is divided between the neutrino and fluid sector, as well as the net lepton number that exited the grid through neutrino emission. After $2.5\,{\rm ms}$, the lepton number lost to matter outflows is negligible. We see that the MC-HR simulation has relatively small violations of lepton number conservation, at the level of $0.5\%$ of the initial number of leptons. The M1-Nulib simulation performs similarly in that respect, and the M1-Radice simulation slightly worse ($3\%$ lepton number violation). This is better than for schemes that do not evolve the number density of neutrinos [38], though still a measurable error over the course of the entire simulation and likely indicative of the fact that inaccuracies created by the joint evolution of the number and energy density of neutrinos occasionally lead to unphysical neutrino energies and corrections applied to the number density (as the M1 schemes are exactly lepton-number conserving up to such corrections). It is maybe surprising that the MC simulation shows better lepton number conservation than most M1 simulations, despite using a scheme that only conserve lepton number up to statistical noise in neutrino-matter interactions. For comparision, errors in the conservation of the total baryon number are $O(10^{-4})$ or less. As already mentioned, the M1-SpEC simulation has large violations of the lepton number conservation, to an unacceptable level in dense regions. Another notable difference between the MC-HR simulation and the M1-NuLib and M1-Radice simulation is that the former has $\sim 50\%$ more $\bar{\nu}_{e}$ neutrinos within the remnant than the latter. All simulations on the other hand have similar numbers of $\nu_{e}$ (and an order of magnitude less $\nu_{e}$ than $\bar{\nu}_{e}$). It is unclear whether such differences can explain the more rapid collapse of the neutron star in MC simulations, or if instead they are a consequence of the more rapid evolution of these simulations towards collapse… or if the larger number of neutrinos in MC is simply due to the fact that the MC code is energy-dependent and thus captures more accurately the evolution of low-energy neutrinos within the remnant.
@@ IV Discussion
Our simulations also allow us to assess more directly the impact of improvements to the M1 scheme recently proposed by Radice et al [39]. We find indications that these improvements do indeed modify the behavior of neutrinos in regions with high scattering optical depths, as predicted in [39], and generally lead to differences with our preexisting M1 code that are often comparable to or larger than other sources of uncertainties in the neutrino transport algorithm. We note in addition that the M1-SpEC simulation has a peculiar behavior leading to the emission of low-energy $\bar{\nu}_{e}$ during merger, violating conservation of electron lepton number. Considering that it occurs in the same high density, cold regions where the M1-SpEC is expected to be least accurate, it is likely a result of the inaccuracies of that M1 scheme. It is also interesting to note that, somewhat surprisingly, even the better behaved M1 simulations (M1-NuLib, M1-Radice) at best conserve lepton number as well as the MC scheme, even though the MC scheme is only constructed to conserve lepton number up to statistical noise in neutrino-matter interactions. The $O(1\%)$ violations of total lepton number conservation observed here are not expected to be significant for the short simulations presented in this manuscript, but may be a worry when applying these methods to much longer numerical evolution without additional care to ensure lepton number conservation.
The M1-Radice and MC simulations show $\sim 20\%$ level agreements in early-time neutrino luminosities, while more significant differences are observed between those three simulations and those using the approximate implicit solve previously implemented in SpEC. Interestingly, this difference between M1-Radice and M1-SpEC also appears sufficient to explain the largest difference between M1 and MC results observed in our previous comparison of low-mass binary neutron star systems, i.e. the fact that the M1 scheme was underproducing $\nu_{x}$ neutrinos by up to $\sim 50\%$. Other differences between M1 and MC remain however notable: the M1 simulations result in overdensities of neutrinos in the polar regions, as was already known, and the MC simulation show higher densities of $\bar{\nu}_{e}$ within the remnant itself. The latter effect is however currently impossible to disentangle from the differing evolution of the MC and M1 simulations towards collapse (and seems equally likely to be a cause or an effect of that rapid collapse, or entirely unrelated to that question).